OpenStudy (anonymous):

If B is between A and C find the value of BC. AB = 3, BC = 5x - 1, AC = 12

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

AC=AB+BC , so u hav 12=3+BC and so BC=9